Suppose time taken to grade an exam-paper by an examiner has mean 12.8 min. and standard deviation 5 min. If the examiner has to grade 46 exam-papers, what is the approximate probability that she will finish grading in less than 609 min. (assuming grading times are independent and she grades continuously)? Answer to 4 decimal place
Answer #1

\sum X ~ N( n \mu , n \sigma 2 ) = (46 * 12.8 , 46 * 52 ) = (588.8 , 1150)

Using central limit theorem ,

(\sum X < x) = P(Z < x -  \mu / sqrt (n \sigma 2 ) ) )

So,

P(\sum X < 609= sqrt ( P(Z < 609 - 588.8 / sqrt(1150) )

= P(Z < 0.60)

= 0.7257